A storm 10.0 km in diameter has wind speeds of 10.0 m/s. What is its angular velocity in radians per second?

a.

3.18 x 10-4 rad/s

b.

6.37 x 10-4 rad/s

c.

1.27 x 10-3 rad/s

d.

2.00 x 10-3 rad/s
Answer #1

Angular velocity (W) =v/r

Where v= velocity

r=radius of storm.

W= 10/5000

Angular velocity (W) =2×10-3 rad/s.